# Zero-shot Context Biasing with Trie-based Decoding using Synthetic Multi-Pronunciation

### v1

@@ Abstract
Contextual automatic speech recognition (ASR) systems allow for recognizing out-of-vocabulary (OOV) words, such as named entities or rare words. However, it remains challenging due to limited training data and ambiguous or inconsistent pronunciations. In this paper, we propose a synthesis-driven multi-pronunciation contextual biasing method that performs zero-shot contextual ASR on a pretrained Whisper model. Specifically, we leverage text-to-speech (TTS) systems to synthesize diverse speech samples containing each target rare word, and then use the pretrained Whisper model to extract multiple predicted pronunciation variants. These variant token sequences are compiled into a prefix-trie, which assigns rewards to beam hypotheses in a shallow-fusion manner during beam-search decoding. After which, any recognized variant is mapped back to the original rare word in the final transcription. The evaluation results on the Librispeech dataset show that our method reduces biased word error rate (WER) by 42% on test-clean and 43% on test-other while maintaining unbiased WER essentially unchanged.
@@ 1 Introduction
In this paper, we introduce a novel approach to enhancing the recognition accuracy of rare words in a pretrained ASR system without requiring any adaptation. Our method leverages synthetic speech to generate diverse pronunciation variants for each hotword, which are then transcribed by the same ASR model. Technically, we synthesize speech using a variety of TTS systems and voices for each templated hotword sample and apply Whisper-large-v3 model to transcribe them. We integrate a multi-pronunciation prefix-trie-based hotword module into the same Whisper-large-v3 model in a shallow-fusion manner, where the prefix-trie is constructed using extracted multiple pronunciation variants of each hotword.
Our experiments on the Librispeech test sets demonstrate that the proposed method reduces WER/B-WER of the baseline Whisper-large-v3 model without affecting the recognition performance of common words. To the best of our knowledge, this is the first work to leverage synthesized multi-pronunciation variants for contextual ASR without requiring model adaptation.
@@ 2.1 Multi-Pronunciation Hotword Speech Synthesis via TTS
To ensure that the hotword appears in a fixed, easily locatable context, we embed it in two minimal utterances: “Start <Hotword> End” and “Begin <Hotword>”. The context can be customized, but we adopt such simple and rigid templates to facilitate reliable extraction from Whisper decoding results and to ensure stable and consistent synthesis from TTS models. Longer or more complex contexts tend to introduce unwanted variability in TTS outputs, which can compromise the reliability of anchor-based hotword localization, as the surrounding context serves as a temporal marker for extracting hotword speech variations. Each engine renders both utterances with three distinct voices and accents from [14], including female British accent, male British accent, and male American accent. This process yields a total of 3 engines × 2 utterances × 3 voices = 18 synthesized utterances per hotword, capturing a rich and comprehensive variety of pronunciation patterns.
@@ 2.2 Multi-Pronunciation Hotword Token Extraction
Figure 2: An illustration of Single-Pron-Hotword and Multi-Pron-Hotword prefix-trie using Final-token-only and Uniform per-token reward schemes. In Single-Pron-Hotword prefix-trie, each hotword corresponds to only one path, standing for the correct way of spelling. In contrast, in a Multi-Pron-Hotword prefix-trie, each hotword may correspond to multiple paths with different spellings. When using the Final reward, only the last state of a valid path will give a reward of 1; whereas when using Uniform reward, each active state will credit a reward of 1 when matched. Numbers in the parentheses stand for the rewards gained till the current state. Red dashed lines are examples of failing to match in the middle state of a prefix-trie path.
Next, each synthesized waveform is fed into the vanilla Whisper-large-v3 model to obtain the recognition results, where we require not only the final sentences but also the sequences of byte-pair encoding (BPE) tokens. This is to align with the requirement of prefix-trie as mentioned in Section 2.3, which performs hotword biasing through BPE-level rewarding scores in each auto-regressive decoding step of the Whisper decoder that generates the probability distribution of the entire BPE dictionary.
@@ 2.3 Prefix-Trie Implementation in Whisper Decoder
We then attempt to advance the hypothesis’ current trie state with $y_{t}$: if $y_{t}$ continues a valid hotword prefix conditioned on $y_{t-1}$, the trie returns a reward $\rho_{y_{t-1}\rightarrow y_{t}}=1$; otherwise we remove the accumulated reward of partial unsuccessful match from the hypothesis’s cumulative score. We define the reward as $\sum_{i=1}^{n}\rho_{y_{t-i}\rightarrow y_{t-i+1}}$ where $n$ is the depth from root to current state of prefix-trie as shown in Figure 2. We form the final score $S$ augmented with contextual rewards, as in Eq. 3, which is then used in place of $S_{\textbf{W}}$} when ranking and pruning beam candidates.
@@
Through this three-stage methodology, as shown in Figure 1, we incorporate a prefix-trie-based multiple pronunciation hotword mechanism with Whisper for contextual biasing, leading to significant reductions in biased-WER without degrading overall ASR accuracy.
@@ 3.1.1 Data and Hotword Selection
From the rare word vocabulary, we select a subset of rare words that exhibit a non-zero initial word-error rate (WER) when decoded with the official Whisper-large-v3 model without contextual biasing. We then perform speech synthesis on this subset. In contrast, rare words that are correctly transcribed are explicitly excluded from further processing and do not go through the synthesis pipeline. Both sections of rare word vocabulary above are taken into account when calculating biased-WER.
Table 1: Statistics of the Librispeech dataset. Appearing rare words refers to the rare words counts that are present in the reference transcripts.
@@
Rare word rates 10.96% 10.22%
@@ 3.2 Experimental Setup
All of our experiments are based on the Whisper-large-v3 model, and we use beam-search decoding with a beam size of 10. We use the vanilla Whisper-large-v3 model as a baseline, comparing against the Whisper-large-v3 with prefix-trie hotword module under various configurations as illustrated in Figure 2:
@@
  • Multi-Pron-Hotword: As proposed in Chapter 2, each hotword is bound with several selected pronunciations generated through a TTS-Whisper pipeline, and may correspond to several paths in the prefix-trie.
@@ 4 Results
Table 2: Performance of proposed multi-pronunciation hotword on Librispeech dataset using biasing list with $N=1000$. SF indicates whether we apply syllable filtering to the multi-pronunciation variants. Reward indicates the reward patterns (Final-token-only vs. Uniform per-token) used. Final-token-only indicates only the final prefix-trie state of a valid path has a reward of 1 while the previous states has a reward of 0. Uniform per-token indicates every prefix-trie state of a valid path has a reward of 1. WER refers to the overall word error rate. B-WER/U-WER refers to biasing/unbiasing WER, indicating the model’s ability to recognize rare/common words.
@@
Baseline Vanilla Whisper-large-v3 - 2.81 1.94 9.86 4.84 3.39 17.55
@@
In Table 2, with the reward scheme set to Final, comparing S1 and the Baseline system, we observe a significant B-WER reduction (19% for test-clean and 22% for test-other), along with slightly better U-WER results. This suggests that a prefix-trie hotword implementation already provides substantial improvements in hotword recognition for Whisper. When we implement the multi-pronunciation hotword strategy without using SF (as in S2 and S3), the U-WER performance degrades dramatically. At the same time, B-WER does not improve at all, showing that our pipeline for generating the multi-pronunciation hotwords brings us a large amount of noisy variations, leading to a wrong rewarding paradigm. However, when SF is applied to remove those faulty variations, comparing S5 with S1, we can conclude that the strategy using multi-pronunciation hotwords with their original forms introduces further B-WER improvement by 10% and 7% for test-clean and test-other, respectively, while the U-WER remains essentially unchanged. Since the rare words only occupy around 10% of the entire test sets, S5 only achieves 2-3% improvement on the overall WER.
Using Uniform reward scheme, comparing S6 with S1, U-WER and B-WER have identical improvement for both test sets, with around 3% performance improvement on U-WER, and 23% B-WER reduction. Also, we observe the same trends when comparing S7 against S5. This indicates that guiding Whisper’s decoding process by rewarding step by step to flavor rare words achieves even better performance. Our final system, S7, integrates a prefix-trie contextual biasing module built on multi-pronunciation and the original form of rare words to the Baseline system, achieving 18%, 21% overall WER reduction, and 43%, 44% B-WER performance improvement, for test-clean and test-other, respectively.
@@ 5 Conclusion
We introduced a novel approach to enhancing zero-shot contextual ASR performance by leveraging synthetic speech for Whisper. Our method involves synthesizing diverse hotword speech via TTS engines and integrating these multi-pronunciation variants extracted from the Whisper model into a prefix-trie for shallow-fusion in beam-search decoding. Without any model fine-tuning or reliance on large language models, our system consistently improves hotword recognition performance on the Librispeech test sets with a hotword biasing list, achieving competitive biased-WER while maintaining and even improving general ASR quality. Our results surpass many fine-tuned contextual ASR systems, making the approach practical, low-cost, and easy to deploy, as it does not require any fine-tuning of the underlying ASR model. Future work will expand the TTS augmentation pipeline by incorporating additional voice styles and more diverse TTS engines, including multilingual and expressive speech synthesis. This would enable coverage of broader pronunciation variability and better simulate real-world user scenarios across regions and accents.

### v2

@@ Abstract
Contextual automatic speech recognition (ASR) systems allow for recognizing out-of-vocabulary (OOV) words, such as named entities or rare words. However, it remains challenging due to limited training data and ambiguous or inconsistent pronunciations. In this paper, we propose a synthesis-driven multi-pronunciation contextual biasing method that performs zero-shot contextual ASR on a pretrained Whisper model. Specifically, we leverage text-to-speech (TTS) systems to synthesize diverse speech samples containing each target rare word, and then use the pretrained Whisper model to extract multiple predicted pronunciation variants. These variant token sequences are compiled into a prefix-trie, which assigns rewards to beam hypotheses in a shallow-fusion manner during beam-search decoding. Subsequently, any recognized variant is mapped back to the original rare word in the final transcription. The evaluation results on the LibriSpeech dataset show that our method reduces biased-word error rate (B-WER) by 43% on test-clean and 44% on test-other while maintaining unbiased-WER (U-WER) essentially unchanged.
@@ 1 Introduction
In this paper, we introduce a novel approach to enhancing the recognition accuracy of rare words in a pretrained ASR system without requiring any adaptation. Our method leverages synthetic speech to generate diverse pronunciation variants for each hotword, which are then transcribed by the same ASR model. Technically, we synthesize speech using a variety of TTS systems and voices for each templated hotword sample and apply Whisper Large-v3 model to transcribe them. We integrate a multi-pronunciation prefix-trie-based hotword module into the same Whisper Large-v3 model in a shallow-fusion manner, where the prefix-trie is constructed using extracted multiple pronunciation variants of each hotword.
Our experiments on the LibriSpeech test sets demonstrate that the proposed method reduces WER/B-WER of the baseline Whisper Large-v3 model without affecting the recognition performance of common words. To the best of our knowledge, this is the first work to leverage synthesized multi-pronunciation variants for contextual ASR without requiring model adaptation.
@@ 2.1 Multi-Pronunciation Hotword Speech Synthesis via TTS
To ensure that the hotword appears in a fixed, easily locatable context, we embed it in two minimal utterances: “Start <Hotword> End” and “Begin <Hotword>”. The context can be customized, but we adopt such simple and rigid templates to facilitate reliable extraction from Whisper decoding results and to ensure stable and consistent synthesis from TTS models. Longer or more complex contexts tend to introduce unwanted variability in TTS outputs, which can compromise the reliability of anchor-based hotword localization, as the surrounding context serves as a temporal marker for extracting hotword speech variations. Each engine renders both utterances with three distinct voices and accents from [14], including British female, British male, and American male. This process yields a total of 3 engines × 2 utterances × 3 voices = 18 synthesized utterances per hotword, capturing a rich and comprehensive variety of pronunciation patterns.
@@ 2.2 Multi-Pronunciation Hotword Token Extraction
Figure 2: An illustration of Single-Pron-Hotword and Multi-Pron-Hotword prefix-trie using Final-token-only and Uniform per-token reward schemes. In Single-Pron-Hotword prefix-trie, each hotword corresponds to only one path, representing the correct spelling. In contrast, in a Multi-Pron-Hotword prefix-trie, each hotword may correspond to multiple paths with different spellings. When using the Final reward, only the last state of a valid path will give a reward of 1; whereas when using Uniform reward, each active state will credit a reward of 1 when matched. Numbers in the parentheses stand for the rewards gained up to the current state. Red dashed lines are examples of failing to match in the middle state of a prefix-trie path.
Next, each synthesized waveform is fed into the vanilla Whisper Large-v3 model to obtain the recognition results, where we require not only the final sentences but also the sequences of byte-pair encoding (BPE) tokens. This is to align with the requirement of prefix-trie as mentioned in Section 2.3, which performs hotword biasing through BPE-level rewarding scores in each auto-regressive decoding step of the Whisper decoder that generates the probability distribution of the entire BPE dictionary.
@@ 2.3 Prefix-Trie Implementation in Whisper Decoder
We then attempt to advance the hypothesis’s current trie state with $y_{t}$: if $y_{t}$ continues a valid hotword prefix conditioned on $y_{t-1}$, the trie returns a reward $\rho_{y_{t-1}\rightarrow y_{t}}=1$; otherwise we remove the accumulated reward of partial unsuccessful match from the hypothesis’s cumulative score. We define the reward as $\sum_{i=1}^{n}\rho_{y_{t-i}\rightarrow y_{t-i+1}}$ where $n$ is the depth from root to current state of prefix-trie as shown in Figure 2. We form the final score $S$ augmented with contextual rewards, as in Eq. 3, which is then used in place of $S_{\textbf{W}}$} when ranking and pruning beam candidates.
@@
Through this three-stage methodology, as shown in Figure 1, we incorporate a prefix-trie-based multiple pronunciation hotword mechanism with Whisper for contextual biasing, leading to significant reductions in B-WER without degrading overall ASR accuracy.
@@ 3.1.1 Data and Hotword Selection
From the rare word vocabulary, we select a subset of rare words that exhibit a non-zero initial WER when decoded with the official Whisper Large-v3 model without contextual biasing. We then perform speech synthesis on this subset. In contrast, rare words that are correctly transcribed are explicitly excluded from further processing and do not go through the synthesis pipeline. Both sections of rare word vocabulary above are taken into account when calculating B-WER.
Table 1: Statistics of the LibriSpeech dataset. Appearing rare words refers to the rare word counts that are present in the reference transcripts.
@@
Rare word rate 10.96% 10.22%
@@ 3.2 Experimental Setup
All of our experiments are based on the Whisper Large-v3 model, and we use beam-search decoding with a beam size of 10. We use the vanilla Whisper Large-v3 model as a baseline, comparing against the Whisper Large-v3 with prefix-trie hotword module under various configurations as illustrated in Figure 2:
@@
  • Multi-Pron-Hotword: As proposed in Section 2, each hotword is bound with several selected pronunciations generated through a TTS-Whisper pipeline, and may correspond to several paths in the prefix-trie.
@@ 4 Results
Table 2: Performance of the proposed multi-pronunciation hotword method on LibriSpeech dataset using biasing list with $N=1000$. SF indicates whether we apply syllable filtering to the multi-pronunciation variants. Reward indicates the reward patterns (Final-token-only vs. Uniform per-token) used. Final-token-only indicates only the final prefix-trie state of a valid path has a reward of 1 while the previous states have a reward of 0. Uniform per-token indicates every prefix-trie state of a valid path has a reward of 1. WER refers to the overall word error rate. B-WER/U-WER refers to biased/unbiased-WER, indicating the model’s ability to recognize rare/common words.
@@
Baseline Vanilla Whisper Large-v3 - 2.81 1.94 9.86 4.84 3.39 17.55
@@
In Table 2, with the reward scheme set to Final, comparing S1 and the Baseline system, we observe a significant B-WER reduction (19% for test-clean and 22% for test-other), along with slightly better U-WER results. This suggests that a prefix-trie hotword implementation already provides substantial improvements in hotword recognition for Whisper. When we implement the multi-pronunciation hotword strategy without using SF (as in S2 and S3), the U-WER performance degrades dramatically. At the same time, B-WER does not improve at all, showing that our pipeline for generating the multi-pronunciation hotwords brings us a large amount of noisy variations, leading to an incorrect rewarding paradigm. However, when SF is applied to remove those faulty variations, comparing S5 with S1, we can conclude that the strategy using multi-pronunciation hotwords with their original forms introduces further B-WER improvement by 10% and 7% for test-clean and test-other, respectively, while the U-WER remains essentially unchanged. Since the rare words only occupy around 10% of the entire test sets, S5 only achieves 2-3% improvement on the overall WER.
Using Uniform reward scheme, comparing S6 with S1, U-WER and B-WER have identical improvement for both test sets, with around 3% performance improvement on U-WER, and 23% B-WER reduction. Also, we observe the same trends when comparing S7 against S5. This indicates that guiding Whisper’s decoding process by rewarding step by step to favor rare words achieves even better performance. Our final system, S7, integrates a prefix-trie contextual biasing module built on multi-pronunciation and the original form of rare words to the Baseline system, achieving 18% and 21% overall WER reduction, as well as 43% and 44% B-WER performance improvement, for test-clean and test-other, respectively.
@@ 5 Conclusion
We introduced a novel approach to enhancing zero-shot contextual ASR performance by leveraging synthetic speech for Whisper. Our method involves synthesizing diverse hotword speech via TTS engines and integrating these multi-pronunciation variants extracted from the Whisper model into a prefix-trie for shallow-fusion in beam-search decoding. Without any model fine-tuning or reliance on large language models, our system consistently improves hotword recognition performance on the LibriSpeech test sets with a hotword biasing list, achieving competitive B-WER while maintaining and even improving general ASR quality. Our results surpass many fine-tuned contextual ASR systems, making the approach practical, low-cost, and easy to deploy, as it does not require any fine-tuning of the underlying ASR model. Future work will expand the TTS augmentation pipeline by incorporating additional voice styles and more diverse TTS engines, including multilingual and expressive speech synthesis. This would enable coverage of broader pronunciation variability and better simulate real-world user scenarios across regions and accents.
Acknowledgment
This research is supported by the RIE2025 Industry Alignment Fund - Industry Collaboration Projects (IAF-ICP) (Award I2301E0026), administered by A*STAR, as well as supported by Alibaba Group and NTU Singapore through Alibaba-NTU Global e-Sustainability CorpLab (ANGEL).